You own 11 compact discs (CDs). You want to randomly arrange 5 of them in a CD rack. What is the probability that the rack ends up in alphabetical order? Be sure to leave your answer as a fraction in order to earn credit.

There is a __________ probability that the rack ends up in alphabetical order.
Answer #1

Since there is only one arrangement out of 9! where the CD's are in alphabetical order, the probability that the rack ends up in alphabetical order is 1/5! = 1/120 = 0.0083